# Trackastra: Transformer-based cell tracking for live-cell microscopy

### v1

@@ 3.2 Bacteria colony tracking (Bacteria)
Here we use one of the largest public bacteria tracking datasets from [51]. This dataset (denoted Bacteria) consists of 39 videos of six different types of bacteria colonies, containing roughly 100K cells with full segmentation and tracking annotations and showing roughly 9K divisions. This dataset is challenging, as objects are densely packed and the colonies grow and divide, leading to large displacements between frames (cf. Fig. 2a). We split it into 31 training, two validation and six test videos, and compare our approach on ground truth detections against various baselines, including Delta 2.0 [34] that was explicitly created for bacterial colony tracking (Tab. 1).
Table 1: Tracking results on Bacteria (using ground truth detections). The test set consists of six videos with a mean of 2912 edges and 236 divisions each. $\downarrow$–metrics report mean absolute errors per video. We show results for three runs per model. The last two rows show results of the general model.
@@
As expected, simple Euclidean-distance-based tracking achieves poor scores, even if a powerful integer linear programming (ILP) linker is used, since the frame rate in this dataset is low and the objects move notably from frame to frame (cf. Fig. 2a, Fig. 3a). A standard procedure for bacteria tracking using the accordingly configured and widely used cell tracking tool TrackMate [46] also leads to poor results, since the movements are often too large even for a generously configured intersection-over-union-based tracker (cf. Fig. 3b). The state-of-the-art deep-learning-based bacteria tracking algorithm Delta 2.0 [34], trained on the same dataset as Trackastra, reduces the total AOGM notably, as shown in Fig. 3c. Surprisingly, training Trackastra using only center point coordinates as features (i.e. without any shape information) performs only slightly worse than Delta 2.0, which in turn does make use of both the grayscale images and the segmentation masks. This highlights the strong cues that object motion alone can provide for biological datasets. When training Trackastra with shallow appearance features (i.e. area and mean intensity of the masks), already using greedy linking based on the predicted associations reduces the total errors (AOGM) by $\sim 70\%$ compared to the state of the art (Tab. 1). When adding an ILP linker, we are further reducing the total error leading to an almost perfect tracking result when compared to Delta 2.0 ($\mathrm{AOGM}=23$ vs. $118$).
@@ 3.3 Nuclei tracking (DeepCell)
Next, we evaluate Trackastra on DynamicNuclearNet [39], the largest publicly available dataset for cell nuclei tracking. This dataset (here denoted DeepCell) contains 130 videos of fluorescently labeled nuclei from five different cell types, containing roughly 600K cells with full segmentation and tracking annotations and showing roughly 2K divisions. Due to the comparatively lower rate of cell divisions, we choose to reduce the impact of divisions in the loss reweighting $W$ by setting $\lambda_{\mathrm{div}}=2$ for this dataset. We adhere to the training-validation-test split as defined in [39]. We train a Trackastra model based on ground truth segmentation, as well as a model based on the segmentations from the Caliban pipeline [39] (cf. Tab. 2). We compare our approach to two recent methods that learn linking costs explicitly [39, 1], and take the scores reported in [39].
@@
Table 3: Out-of-domain results on Hela. Specialized models trained only on DeepCell (using ground truth detections), Trackastra-general being trained on a diverse dataset. The test set contains two videos with a mean of 16835 edges and 151 divisions.
@@
To probe the out-of-domain capabilities of Trackastra, we take a model trained on DeepCell and apply it to a similar dataset of fluorescently tagged nuclei (HeLa cells) from the Cell Tracking Challenge (denoted Hela), refer to Tab. 3. Interestingly, Trackastra significantly outperforms Caliban ($\mathrm{AOGM}=190$ vs. $931$), potentially because of its use of only shallow input features (such as positions and basic shape descriptors), which likely helps the model prevent overfitting. Notably, the general model (Trackastra-general, cf. Sec. 3.5) trained on a diverse dataset performs substantially better than the specialized models.
@@ 3.4 ISBI particle tracking challenge
Table 4: Tracking results on the Vesicle dataset from the ISBI particle tracking challenge (using ground truth detections). For a definition of the metrics see [5].
@@
Here we assess how well Trackastra performs for single molecule tracking, a domain closely related to cell tracking, but objects dot not divide. Specifically, we use the challenging Vesicle dataset from the ISBI particle tracking challenge [5] (cf.Fig. 2c). This gives us the chance to compare Trackastra against a state-of-the-art transformer-based tracking approach, MoTT [53], as well as classical methods such as the Kalman filter (KF) [20] and the linear assignment problem (LAP) [17, 8] (cf. Tab. 4). Interestingly, even in this context Trackastra outperforms MoTT for the case of high vesicle density when using ground truth detections in terms of the association accuracy ($\alpha,\beta$) as well as Jaccard similarity coefficient ($JSC_{\theta}$), which is a measure of the overlap between the predicted and ground truth tracks (Tab. 4). This highlights the general applicability of our approach, as no adjustments to the model or the training procedure were required.
@@ 3.5 Multi-domain general Trackastra model
We train a Trackastra model on a diverse dataset, including all data used in the experiments described before (Secs. 3.2, 3.3 and 3.4), as well as additional data from the Cell Tracking Challenge and from [21, 40]. We evaluate this model (denoted Trackastra-general) on the Bacteria and DeepCell test sets. Notably, on both Bacteria and DeepCell, the general model outperforms the specialized single-domain Trackastra models (cf. Tabs. 1 and 2). Furthermore, when applying the general model to the out-of-domain Hela dataset (which was not part of any training set), it outperforms the specialised Trackastra model trained on DeepCell, demonstrating the importance of a large, diverse training set for out-of-domain tracking performance (cf. Tab. 3).
@@ 3.6 Ablations
Table 5: Ablation for parental softmax on Bacteria (using ground truth detections). The test set consists of six videos with a mean of 2912 edges and 236 divisions. Lower is better. We report the results for two runs per model.
Linking Parental softmax AOGM FP edges FN edges FP divs FN divs
@@
Figure 4: Ablations on Bacteria (using ground truth detections) using only center points as features, and a LAP linker [46]. Lower is better. We show results for three runs per model.
@@
Here we vary the number $L$ of attention layers in both the encoder and decoder, as well as the embedding dimension $d$, using slightly smaller Trackastra models for computational efficiency (cf. Figs. 4(a) and 4(b)). As expected, using no attention layers ($L=0$), i.e. directly predicting the association matrix with the two projection heads leads to poor performance, confirming the importance of attention across all detections. Furthermore, using more than $L=6$ layers does not lead to a notable improvement.
@@
We run experiments with different window sizes $s$, i.e. different temporal context available to the model (cf. Fig. 4(c)). When using only $s=2$ we find that the performance substantially deteriorates, whereas window sizes $s\in(3,6)$ all lead to good results, demonstrating that a relatively small temporal context is already enough for decent tracking results.
Table 6: Ablations on Bacteria (using ground truth detections). The test set consists of six videos with a mean of 2912 edges and 236 divisions each. We show results for three runs per model.
@@
Finally, we ablate some basic components of Trackastra in Tab. 6. Removing positional encodings as well as the basic shape descriptors from Trackastra increases errors notably, while the decrease in performance when replacing the tailored ILP optimizer with a simple greedy linker is less pronounced.
@@ 4 Discussion
We presented Trackastra, a robust method to track dividing objects such as cells with a powerful and scalable transformer architecture. Trackastra uses only object positions and shallow object features as input, which makes it readily applicable to new scenarios whenever a domain specific detection method is already available. We demonstrate that Trackastra works well for different imaging modalities and biological model systems that share the particularity of dividing, yet non-fusing objects. Furthermore, we show that a single model is able to generalize well to datasets from various different domains, which is crucial for practitioners. An important limitation of the presented model is that it cannot correct faulty detection inputs, but we envision that training a detection model and Trackastra end-to-end will be beneficial. Furthermore, we currently don’t make use of the pair-wise association predictions from non-adjacent time frames, which we expect to be beneficial for circumventing faulty detections, for example for track gap closing in case of missing detections. Finally, while the presented results are limited to 2d datasets, Trackastra promises to scale well to 3d datasets as the architecture does not process dense images. Overall our work demonstrates the potential that pure transformer-based architectures can hold for the field of cell tracking.
4.0.1 Acknowledgements
We thank Arlo Sheridan, Talmo Pereira, Uwe Schmidt and Albert Dominguez for helpful discussions, Simon van Vliet and colleagues for creating the large-scale annotated Bacteria dataset, and the EPFL School of Life Sciences ELISIR program and CARIGEST SA for their generous funding.

### v2

@@ 3.2 Bacteria colony tracking (Bacteria)
Here we use one of the largest public bacteria tracking datasets [55]. This dataset (denoted Bacteria) consists of 39 videos of six different types of bacteria colonies, containing roughly 100k cells with full segmentation and tracking annotations and showing roughly 9k divisions. This dataset is challenging, as objects are densely packed and the colonies grow and divide, leading to large displacements between frames (cf. Fig. 2a). We split it into 31 training, two validation and six test videos, and compare our approach on ground truth detections against various baselines, including Delta 2.0 [37] that was explicitly created for bacterial colony tracking (Tab. 1).
Table 1: Tracking results on Bacteria (using ground truth detections). The test set consists of six videos with a mean of 2912 edges and 236 divisions. $\downarrow$–metrics report mean absolute errors per video. We show results for three runs per model. The last two rows show results of the general model (cf. Sec. 3.5).
@@
As expected, simple Euclidean-distance-based tracking achieves poor scores, even if a powerful integer linear programming (ILP) linker is used, since the frame rate in this dataset is low and the objects move notably from frame to frame (cf. Fig. 2a, Fig. 3a). A standard procedure for bacteria tracking using the accordingly configured and widely used cell tracking tool TrackMate [50] also leads to poor results, since the movements are often too large even for a generously configured intersection-over-union-based tracker (cf. Fig. 3b). The state-of-the-art deep-learning-based bacteria tracking algorithm Delta 2.0 [37], trained on the same dataset as Trackastra, reduces the total AOGM notably, as shown in Fig. 3c. Surprisingly, training Trackastra using only center point coordinates as features (i.e. without any appearance information) performs only slightly worse than Delta 2.0, which in turn does make use of both the grayscale images and the segmentation masks. This highlights the strong cues that object motion alone can provide for biological datasets. When training Trackastra with shallow object features, already using greedy linking based on the predicted associations reduces the total errors (AOGM) by $\sim 70\%$ compared to the state of the art (Tab. 1, Fig. 3d, Vid. S1). When adding an ILP linker, we are further reducing the total error to an almost perfect tracking result ($\mathrm{AOGM}=23$ vs. $118$ for Delta 2.0).
@@ 3.3 Nuclei tracking (DeepCell)
Next, we evaluate Trackastra on DynamicNuclearNet [42], the largest publicly available dataset for cell nuclei tracking. This dataset (here denoted DeepCell) contains 130 videos of fluorescently labeled nuclei from five different cell types, containing roughly 600k cells with full segmentation and tracking annotations and showing roughly 2k divisions. Due to the comparatively lower rate of cell divisions, we choose to reduce the impact of divisions in the loss reweighting $W$ by setting $\lambda_{\mathrm{div}}=2$ for this dataset. We adhere to the training-validation-test split as defined in [42]. We train a Trackastra model based on ground truth segmentations, as well as a model based on the segmentations from the Caliban pipeline [42] (cf. Tab. 2). We compare our approach to two recent methods that learn linking costs explicitly [42, 1], and take the scores reported in [42].
@@
Table 3: Out-of-domain results on Hela. Specialized models trained only on DeepCell (using ground truth detections), Trackastra-general trained on a diverse dataset. The test set contains two videos with a mean of 16835 edges and 151 divisions.
@@
To probe the out-of-domain capabilities of Trackastra, we take a model trained on DeepCell and apply it to a similar dataset of fluorescently tagged nuclei of HeLa cells from the Cell Tracking Challenge (denoted Hela), refer to Tab. 3. Interestingly, Trackastra significantly outperforms Caliban ($\mathrm{AOGM}=190$ vs. $931$), potentially because of its use of only shallow input features (such as positions and basic shape descriptors), which likely helps the model prevent overfitting. Notably, the general model (Trackastra-general, cf. Sec. 3.5) trained on a diverse dataset performs substantially better than the specialized models.
@@ 3.4 ISBI particle tracking challenge
Table 4: Tracking results on the Vesicle dataset from the ISBI particle tracking challenge (using GT detections). For a definition of the metrics see [5].
@@
Here we assess how well Trackastra performs for single molecule tracking, a domain closely related to cell tracking where, however, objects do not divide. Specifically, we use the challenging Vesicle dataset from the ISBI particle tracking challenge [5] (cf. Fig. 2c). This gives us the chance to compare Trackastra against a state-of-the-art transformer-based tracking approach, MoTT [58], as well as classical methods such as the Kalman filter (KF) [22] and the linear assignment problem (LAP) [19, 8] (cf. Tab. 4). Interestingly, even in this context Trackastra outperforms MoTT for the case of high vesicle density when using ground truth detections in terms of the association accuracy ($\alpha,\beta$) as well as Jaccard similarity coefficient ($JSC_{\theta}$), which is a measure of the overlap between the predicted and ground truth tracks (Tab. 4). This highlights the general applicability of our approach, as no adjustments to the model or the training procedure were required.
@@ 3.5 Multi-domain general Trackastra model
We train a Trackastra model on a diverse dataset, including all data used in the experiments described before (Secs. 3.2, 3.3 and 3.4), as well as additional data from the Cell Tracking Challenge [34] and from [23, 43, 56, 11]. We evaluate this model (denoted Trackastra-general) on the Bacteria and DeepCell test sets. Notably, on both Bacteria and DeepCell, the general model outperforms the specialized single-domain Trackastra models (cf. Tabs. 1 and 2). Furthermore, when applying the general model to the out-of-domain Hela dataset (which was not part of any training set), it outperforms the specialized Trackastra model trained on DeepCell, demonstrating the importance of a large, diverse training set for out-of-domain tracking performance (cf. Tab. 3).
Finally, we also submit a single Trackastra-general model to the cell linking benchmark of the Cell Tracking Challenge (7th edition, 2024). Our model performs best overall according to structural as well as biological metrics across 13 diverse datasets, as well as best individually for roughly half of the datasets [33].
@@ 3.6 Ablations
Table 5: Ablation for parental softmax on Bacteria (using GT detections). Test set as in Tab. 1. We report results for two runs per model.
Linking Parental softmax AOGM\>$\downarrow$ FP edges\>$\downarrow$ FN edges\>$\downarrow$ FP divs\>$\downarrow$ FN divs\>$\downarrow$
@@
Figure 4: Ablations on Bacteria (using ground truth detections) using only center points as features, and a LAP linker. Lower is better. We show results for three runs per model.
@@
Here we vary the number $L$ of attention layers in both the encoder and decoder, as well as the embedding dimension $d$, using slightly smaller Trackastra models for computational efficiency (cf. Fig. 4a,b). As expected, using no attention layers ($L=0$), i.e. directly predicting the association matrix with the two projection heads leads to poor performance, confirming the importance of attention across all detections. Furthermore, using more than $L=6$ layers does not lead to a notable improvement.
@@
We run experiments with different window sizes $s$, i.e. different temporal context available to the model (cf. Fig. 4c). When using only $s=2$ we find that the performance substantially deteriorates, whereas window sizes $s\in(3,6)$ all lead to good results, demonstrating that a relatively small temporal context is already enough for decent tracking results.
Table 6: Ablations on Bacteria (using GT detections). Test set as in Tab. 1. We show results for three runs per model.
@@
Finally, we ablate basic components of Trackastra in Tab. 6. Removing positional encodings as well as the basic shape descriptors from Trackastra increases errors notably, while the decrease in performance when replacing the tailored ILP optimizer with a simple greedy linker is less pronounced.
@@ 4 Discussion
We presented Trackastra, a robust method to track dividing objects such as cells using a powerful and scalable transformer architecture. Trackastra uses only object positions and shallow object features as input, making it readily applicable to new scenarios whenever a domain specific detection method is already available. We demonstrate that Trackastra performs well across different imaging modalities and biological model systems that share the particularity of dividing, yet non-fusing objects. Additionally, we show that a single model trained on a large cross-modality dataset is able to generalize well to datasets from various different domains, which is crucial for practitioners. Finally, Trackastra has the potential for tracking other types of dividing objects, such as icebergs in satellite imagery. An important limitation of the presented approach is that it currently does not correct faulty detection inputs, but we anticipate that training a detection model and Trackastra end-to-end will address this issue. Furthermore, we currently do not make use of the pairwise association predictions from non-adjacent time frames, which we expect to be beneficial for circumventing faulty detections, for example for track gap closing in case of missing detections. Moreover, while the presented results are limited to 2D datasets, Trackastra is expected to scale well to 3D datasets since the architecture does not require to process dense images. Overall our work demonstrates the potential that pure transformer-based architectures can hold for the field of cell tracking.
Acknowledgements
We thank Arlo Sheridan, Talmo Pereira, Uwe Schmidt and Albert Dominguez for helpful discussions, Morgan Schwartz for releasing the code for the Caliban benchmarks, Simon van Vliet and Johannes Seiffarth for creating large-scale annotated datasets, and the EPFL School of Life Sciences ELISIR program and CARIGEST SA for their generous funding. Additionally we would like to acknowledge the Janelia Trackathon 2023 and the resulting metrics library traccuracy.